OpenStudy (anonymous):

X+2y=2 in slope intercept form

OpenStudy (anonymous):

slope intercept form: y=mx+b b is the constant

OpenStudy (anonymous):

Can you explain it to me? eLg

OpenStudy (anonymous):

The web site answers any question well almost any questoin

jimthompson5910 (jim_thompson5910):

isolate y to get \[\large x+2y=2\] \[\large 2y=2-x\] \[\large 2y=-x+2\] \[\large y=\frac{-x+2}{2}\] \[\large y=\frac{-x}{2}+\frac{2}{2}\] \[\large y=-\frac{1}{2}x+1\] Hopefully these steps will help you solve other problems like this

OpenStudy (anonymous):

jim_thompson5910 showed the steps... you can ask me if you did not understand it.